# Curvature pinching of asymptotically conical gradient expanding Ricci solitons

### v2

@@ Abstract.
Since the well-known work of Hamilton [62] and Ivey [64], the Hamilton-Ivey curvature pinching and its generalizations have become a signature feature of gradient shrinking and steady Ricci solitons, and more generally, of ancient solutions to the Ricci flow. However, analogous results for gradient expanding Ricci solitons have remained elusive. This is largely due to the fact that the proofs of existing curvature pinching estimates crucially rely on shrinking and steady solitons being ancient, a property not shared by gradient Ricci expanders. In this paper, we investigate curvature pinching phenomena in non-compact asymptotically conical gradient expanding Ricci solitons and establish several Hamilton-Ivey type curvature pinching estimates. These results are parallel to those known for shrinking and steady Ricci solitons. In particular, we prove a three-dimensional Hamilton-Ivey type curvature pinching theorem: any three-dimensional non-compact gradient Ricci expander, which is asymptotic to a cone with positive scalar curvature, must have positive sectional curvature. As an application, we combine our result with that of Deruelle [51] to establish a uniqueness theorem for three-dimensional asymptotically conical expanders with positive scalar curvature.
Furthermore, we formulate a general proof method and apply it to obtain analogues of several additional known generalized Hamilton-Ivey type curvature pinching results for ancient solutions. Among these is a curvature pinching estimate for four-dimensional asymptotically conical Ricci expanders with uniformly positive isotropic curvature, analogous to a result for four-dimensional gradient steady solitons due to Brendle [7].
@@ 1. Introduction
The first examples of gradient expanding Ricci solitons are the one-parameter family of complete, rotationally symmetric, asymptotically conical gradient expanding Ricci solitons on $\mathbb{R}^{n}$ ($n\geq 3$) with positive or negative sectional curvature constructed by Bryant [11], and the one-parameter family of complete, $U(n)$-invariant, asymptotically conical gradient expanding Kähler-Ricci solitons on $\mathbb{C}^{n}$ with similar curvature behavior constructed by the first author [16]. The constructions in [15, 16] were later extended by Feldman-Ilmanen-Knopf [56] to produce gradient expanding Kähler-Ricci solitons on the complex line bundles O(k)O(-k) ($k>n$) over the complex projective space ${\mathbb{C}}P^{n}$ ($n\geq 1$), and further generalized by Dancer-Wang [50]. Additional constructions can be found in [65, 2, 49, 57, 12, 85, 86, 77, 41, 42].
@@
Despite the progress described above, a key feature well known for gradient shrinking and steady Ricci solitons has remained absent in the expanding case: Hamilton-Ivey type curvature pinching. As is well-known, a hallmark of the three-dimensional Ricci flow is the Hamilton-Ivey curvature pinching [62, 64] (see also [27, Theorem 2.4.1]), which asserts that when curvature blows up, the positive part blows up at a faster rate than the absolute value of the negative part. In particular, 3-dimensional singularity models that are shrinking or steady gradient Ricci solitons, or more generally ancient solutions, must have nonnegative sectional curvature. This remarkable feature is powerful: it enables the application of the Li-Yau-Hamilton differential Harnack inequality and special geometry of nonnegatively curved 3-manifolds to effectively analyze three-dimensional singularity models. The Hamilton-Ivey curvature pinching was later extended by B.-L. Chen [35] to arbitrary 3-dimensional ancient solutions [^1] Ancient solutions exist for all past time, up to a final singular time; immortal solutions, starting at some initial time, exist for all future time., which include gradient shrinking and steady solitons as important special cases. It implies that any complete ancient solution in dimension three must have nonnegative sectional curvature. This has played a crucial role in the classifications of 3-dimensional gradient shrinking and steady Ricci solitons [20, 6, 66, 67, 68], as well as 3D ancient solutions [8, 9].
@@
In higher dimensions, various forms of the generalized Hamilton-Ivey curvature pinching have been established for ancient solutions under suitable assumptions, see, e.g., [89, 3, 70, 43]. More recently, the authors proved certain curvature pinching properties for $4$-dimensional ancient solutions with positive isotropic curvature (PIC) or half PIC [24, 25], leading to partial classifications of $4$-dimensional shrinking and steady gradient Ricci solitons with weakly positive isotropic curvature (WPIC) or half WPIC.
@@
However, analogous curvature pinching results for gradient expanding Ricci solitons have remained largely unknown. This is partly because the proofs of Hamilton-Ivey-type estimates for shrinking and steady solitons rely in an essential way on the ancient nature of these solutions. Since gradient expanders are not ancient solutions but special immortal solutions, those methods do not apply.
@@
Motivated in part by the second author’s work [87] on the convexity of mean convex, asymptotically conical self-expanders in mean curvature flow, and in part by the close resemblance between curvature estimates for Ricci expanders with nonnegative curvature [22, 23] and those for shrinking or steady Ricci solitons [74, 75, 21, 30, 17] (see especially the comparison in dimension four given in [17]), we began to investigate whether analogous Hamilton-Ivey-type curvature pinching properties might hold in the expanding case.
By adapting an argument from the second author’s recent work [87] on asymptotically conical mean curvature expanders – inspired in turn by Spruck-Xiao [84] and Xie-Yu [88] – we have indeed established a number of generalized Hamilton-Ivey-type curvature pinching results for non-compact, asymptotically conical, gradient expanding Ricci solitons with positive scalar curvature.
@@
Our first result is a Hamilton-Ivey type curvature pinching for $3$-dimensional non-compact, asymptotically conical, gradient expanders with positive scalar curvature.
@@
Let $(M^{3},g,f)$ be a $3$-dimensional non-compact, asymptotically conical gradient expanding Ricci soliton. Suppose the asymptotic cone has positive scalar curvature. Then, $(M^{3},g,f)$ must have positive sectional curvature.
@@
Theorem 1.1 may be viewed as an analogue of B.-L. Chen’s result for three-dimensional ancient solutions to the Ricci flow [35]. Moreover, after our paper was completed, the authors learned from P.-Y. Chan that a similar result to Theorem 1.1 was proved in [34, Corollary 1.16] by a different method.
As an application of Theorem 1.1, for $n=3$, we can extend Deruelle’s uniqueness result [51, Theorem 1.3] to asymptotically conical expanding Ricci solitons with positive scalar curvature.
Corollary 1.1.
Let $(\Sigma^{2},\bar{g})$ be a smooth 2-sphere with Gaussian curvature $\bar{K}>1$. Then, there exists a unique $3$-dimensional gradient expanding Ricci soliton with positive scalar curvature asymptotic to the cone $\mathcal{C}^{3}:=[0,\infty)\times\Sigma^{2}$ with link $\Sigma^{2}$.
For $n\geq 4$, we have the following Hamilton-Ivey type curvature pinching result for asymptotically conical, gradient expanding solitons with positive scalar curvature and vanishing Weyl tensor.
@@
Theorem 1.2 is an analogue of Z.-H. Zhang’s result for locally conformally flat gradient shrinking (and steady) solitons [89]. Moreover, by the work of [19, 26], the locally conformally flat assumption can be replaced by the weaker assumption of vanishing $D$-tensor introduced in [18, 19].
As a consequence, by combining with the work of Cao-Chen [19] and Chen-Wang [39], we have the following application in dimension four.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact, half conformally flat, asymptotically conical gradient expanding Ricci soliton. Suppose the asymptotic cone has positive scalar curvature. Then, $(M^{4},g,f)$ has positive curvature operator.
Our next two results concern curvature pinching of 4-dimensional asymptotically conical gradient expanding Ricci solitons with either positive isotropic curvature (PIC) or half positive isotropic curvature (half-PIC); see Section 2.2 for definitions.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with half positive isotropic curvature. If the asymptotic cone has positive scalar curvature and satisfies either $A\geq 0$ or $C\geq 0$, then $(M^{4},g,f)$ has either $A>0$ or $C>0$.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with positive isotropic curvature. If the asymptotic cone has positive scalar curvature, then the Ricci curvature of $(M^{4},g,f)$ is $2$-positive.
@@
Theorems 1.3 and 1.4 form an analogue of the authors’ previous results for $4$-dimensional complete ancient solutions with half PIC and PIC [24, 25], respectively. Moreover, both results are valid under some slightly weaker assumptions; see Theorem 4.1 and Theorem 4.2 in Section 4.
By observing common patterns in the proofs of Theorem 1.3 and Theorem 1.4, we formulate a general method of proof (Lemma 5.1) and apply it to obtain analogues of several additional known generalized Hamilton-Ivey type curvature pinching results for ancient solutions. This includes the following result for $4$-dimensional asymptotically conical gradient expanding Ricci soliton with uniformly positive isotropic curvature (UPIC), as well as several others stated in Theorem 5.1.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact gradient expanding Ricci soliton with uniformly positive isotropic curvature. If $(M^{4},g,f)$ is asymptotic to a non-flat Euclidean cone, then it is rotationally symmetric.
@@
Acknowledgements. The first author was partially supported by a Simons Fellowship and a grant from the Simons Foundation. The second author would like to thank Prof. Xiaochun Rong for his continual support and encouragement. Both authors would like to thank Dr. P.-Y. Chan for very helpful information that allowed the removal of the positive scalar curvature assumption on $(M^{n},f,g)$ in Theorem 1.1 and Theorem 1.2 in an earlier version of the paper.
@@ 2.2. Curvature decomposition and isotropic curvature of four-manifolds
${Rm}=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\tfrac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\tfrac{R}{12}I\end{pmatrix}$,
@@
Let $A_{1}\leq A_{2}\leq A_{3}$ and $C_{1}\leq C_{2}\leq C_{3}$ denote the eigenvalues of $A$ and $C$, respectively, and let $a_{1}\leq a_{2}\leq a_{3}$ and $c_{1}\leq c_{2}\leq c_{3}$ denote the eigenvalues of $W^{+}$ and $W^{-}$, respectively. Then
$A_{i}=a_{i}+\tfrac{R}{12},\qquad C_{i}=c_{i}+\tfrac{R}{12},\qquad(i=1,2,3).$
Since $a_{1}+a_{2}+a_{3}=\operatorname{tr}W^{+}=0=\operatorname{tr}W^{-}=c_{1}+c_{2}+c_{3}$, it follows that
$\operatorname{tr}A=\operatorname{tr}C=\tfrac{R}{4}.$
@@
It turns out that, in dimension four, these curvature conditions (and their natural extensions, half PIC or half WPIC) can be characterized in terms of the $3\times 3$ matrices $A$ and $C$ as follows:
@@
  • half PIC (half WPIC) if and only if either $A$ or $C$ is 2-positive (weakly 2-positive), i.e., $A_{1}+A_{2}>0$ ($A_{1}+A_{2}\geq 0$) or $C_{1}+C_{2}>0$ ($C_{1}+C_{2}\geq 0$).
@@ 2.3. Basic differential equations satisfied by curvatures of Ricci solitons
Except for the first identity in Lemma 2.1, the factor 22 differs from [60] due to our normalization of the inner product on $\wedge^{2}(M)$ (see [24, (2.4)]). Moreover, the first two equations are valid in all dimensions.
@@ 3. Curvature pinching for asymptotically conical Ricci expanders with vanishing Weyl tensor
This section is devoted to the proof of Theorem 1.1 and Theorem 1.2. For the reader’s convenience, we restate Theorem 1.2 in the following form.
@@
First of all, since the asymptotic cone of $(M^{n},g,f)$ has positive scalar curvature, it follows from the strong maximum principle and [31, Theorem 1.6] that $(M^{n},g,f)$ itself has positive scalar curvature $R>0$. This fact will be used in the proofs of Theorem 3.1 and Theorem 1.1.
Next, shall need the following differential inequality for the ratio between the smallest Ricci eigenvalue and the scalar curvature.
@@
Fix any point $p_{0}\in M$, and choose a unit vector field $e_{1}$ such that $Rc(e_{1})=\lambda_{1}e_{1}$ at $p_{0}$, where $\lambda_{1}$ denotes the smallest Ricci eigenvalue. Extend $e_{1}$ by parallel transport along geodesics emanating from $p_{0}$. Clearly, $\lambda_{1}\leq Rc(e_{1},e_{1})$ with equality at $p_{0}$. Evaluating at $p_{0}$, we obtain, in the barrier sense (cf. Calabi [13]),
@@
where in the last equality, we have used [54, (2.8)] (see also [80, Lemma 2.5]).
@@
Finally, it is clear from the above expression that $h\leq 0$. ∎
Now, we divide the proof of Theorem 3.1 into two parts.
Part I: $(M^{n},g,f)$ has nonnegativity Ricci curvature $Rc\geq 0$.
@@
Case 1: Interior infimum. Suppose the negative infimum $\epsilon_{1}$ is attained at some point $p_{0}\in M$. By Lemma 3.1 and Calabi’s barrier strong maximum principle (Lemma 2.2), the ratio $\lambda_{1}/R$ must be constant on some bounded connected domain $p_{0}\in\Omega\subset M^{-}$. In particular, $h\equiv 0$ on $\Omega$ which forces
@@
On the other hand, since the expanding Ricci soliton satisfies $W\equiv 0$, its asymptotic cone $\mathcal{C}$ also has vanishing Weyl tensor. By Lemma 6.2, $\mathcal{C}$ has nonnegative Ricci curvature, whose smallest eigenvalue $\tilde{\lambda}_{1}\equiv 0$ occurs in the radial direction. Hence
@@
Part II: $(M^{n},g,f)$ has positive curvature operator $Rm>0$.
First of all, when the Weyl tensor $W$ vanishes, we may express the smallest sectional curvature as (see, e.g., [89, Proposition 3.1])
@@
We also note that, for any manifold with vanishing Weyl tensor, $\mu$ also coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [89, Proposition 3.1(i)]). In particular, when $W=0$, nonnegative sectional curvature is equivalent to nonnegative curvature operator $Rm\geq 0$.
@@
Evaluating at $p_{0}$ in the barrier sense, using (3.6) and [54, (2.8)] (see also [80, Lemma 2.5]), we obtain
@@
$\begin{split}\bigl((n-2)\lambda_{2}-R\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}&=-\Bigg(\lambda_{1}+\lambda_{2}+\sum_{j=3}^{n}(\lambda_{j}-\lambda_{2})\Bigg)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &=-(\lambda_{1}+\lambda_{2})\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &\quad-\Biggl(\sum_{k=3}^{n}(\lambda_{k}-\lambda_{2})\Biggr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\end{split}$ (3.8)
@@
Now, we are ready to complete the proof of Theorem 3.1, i.e., $(M^{n},g,f)$ has positive curvature operator $Rm>0$.
Proof of $Rm>0$.
@@
Case 1: Interior infimum. Suppose the negative infimum $\epsilon_{0}$ is attained at some point $p_{0}\in M$. By Theorem 3.1 (i), we know that the Ricci curvature is nonnegative. Then, by Lemma 3.2, we have

### v3

@@ Abstract.
In this paper, we investigate curvature pinching phenomena in complete non-compact asymptotically conical gradient expanding Ricci solitons and establish several Hamilton–Ivey type curvature pinching estimates. These results are parallel to those known for shrinking and steady Ricci solitons. In particular, we prove a three-dimensional Hamilton–Ivey type curvature pinching theorem: any three-dimensional non-compact gradient Ricci expander, which is asymptotic to a cone with positive scalar curvature, must have positive sectional curvature. Furthermore, we formulate a general method and apply it to obtain analogues of several additional known generalized Hamilton–Ivey type curvature pinching results for ancient solutions. Among these is a curvature pinching estimate for four-dimensional asymptotically conical Ricci expanders with uniformly positive isotropic curvature, analogous to a result for four-dimensional gradient steady solitons due to Brendle [8].
@@ 1. Introduction
The first examples of gradient expanding Ricci solitons are the one-parameter family of complete, rotationally symmetric, asymptotically conical gradient expanding Ricci solitons on $\mathbb{R}^{n}$ ($n\geq 3$) with positive (and negative) sectional curvature constructed by Bryant [12], and the one-parameter family of complete, $U(n)$-invariant, asymptotically conical gradient expanding Kähler-Ricci solitons on $\mathbb{C}^{n}$ with similar curvature behavior constructed by the first author [17]. The constructions in [16, 17] were later extended by Feldman-Ilmanen-Knopf [57] to produce gradient expanding Kähler-Ricci solitons on the complex line bundles O(k)O(-k) ($k>n$) over the complex projective space ${\mathbb{C}}P^{n}$ ($n\geq 1$), and further generalized by Dancer-Wang [51]. Additional constructions can be found in [66, 2, 50, 58, 13, 86, 87, 78, 43].
@@
Despite the progress described above, a key feature well known for gradient shrinking and steady Ricci solitons has remained absent in the expanding case: Hamilton–Ivey type curvature pinching. As is well-known, a hallmark of the three-dimensional Ricci flow is the Hamilton–Ivey curvature pinching [63, 65] (see also [29, Theorem 2.4.1]), which asserts that when curvature blows up, the positive part blows up at a faster rate than the absolute value of the negative part. In particular, 3-dimensional singularity models that are shrinking or steady gradient Ricci solitons, or more generally ancient solutions, must have nonnegative sectional curvature. This feature is remarkable: it enables the application of the powerful Li–Yau–Hamilton differential Harnack inequality and the geometry of nonnegatively curved 3-manifolds to effectively analyze three-dimensional singularity models.
The Hamilton–Ivey curvature pinching was later extended by B.-L. Chen [37] to arbitrary 3-dimensional ancient solutions [^1] Ancient solutions exist for all past time, up to a final singular time; immortal solutions, starting at some initial time, exist for all future time., which include gradient shrinking and steady solitons as important special cases. It implies that any complete ancient solution in dimension three must have nonnegative sectional curvature. This has played a crucial role in the classifications of 3-dimensional gradient shrinking and steady Ricci solitons [22, 7, 67, 68, 69], as well as 3D ancient solutions [9, 10].
@@
In higher dimensions, various forms of the generalized Hamilton–Ivey curvature pinching have been established for ancient solutions under suitable assumptions, see, e.g., [90, 3, 71, 44]. More recently, the authors proved certain curvature pinching properties for $4$-dimensional ancient solutions with positive isotropic curvature (PIC) or half-PIC [26, 27], leading to partial classifications of $4$-dimensional shrinking and steady gradient Ricci solitons with weakly positive isotropic curvature (WPIC) or half-WPIC.
@@
However, analogous curvature pinching results for gradient expanding Ricci solitons have remained largely unknown. This is partly because the proofs of Hamilton–Ivey-type estimates for shrinking and steady solitons rely in an essential way on the ancient nature of these solutions. Since gradient expanders are not ancient solutions but special immortal solutions, those methods do not apply.
@@
Motivated in part by the second author’s work [88] on the convexity of mean convex, asymptotically conical self-expanders in mean curvature flow, and in part by the close resemblance between curvature estimates for Ricci expanders with nonnegative curvature [24, 25] and those for shrinking or steady Ricci solitons [75, 76, 23, 32, 18] (see especially the comparison in dimension four given in [18]), we began to investigate whether analogous Hamilton–Ivey-type curvature pinching properties might hold in the expanding case. By adapting an argument from the second author’s recent work [88] on asymptotically conical mean curvature expanders–inspired in turn by Spruck-Xiao [85] and Xie-Yu [89]–we have indeed established a number of generalized Hamilton–Ivey-type curvature pinching results for non-compact, asymptotically conical, gradient expanding Ricci solitons with positive scalar curvature.
@@
Our first result is a Hamilton–Ivey type curvature pinching for $3$-dimensional complete non-compact, asymptotically conical, gradient Ricci expanders with positive scalar curvature.
@@
Let $(M^{3},g,f)$ be a $3$-dimensional non-compact, asymptotically conical [^2] For all our results in this paper, except Corollary 1.2, it suffices to assume $C^{2}$-asymptotics in the sense of Definition 2.1(a). gradient expanding Ricci soliton. Suppose the asymptotic cone has positive scalar curvature. Then, $(M^{3},g,f)$ must have positive sectional curvature.
@@
Theorem 1.1 may be viewed as an analogue of B.-L. Chen’s result for three-dimensional ancient solutions to the Ricci flow [37]. Moreover, after our paper was completed, we learned from P.-Y. Chan that a similar result to Theorem 1.1 was proved in [36, Corollary 1.16] by a different method.
For $n\geq 4$, we have the following Hamilton–Ivey type curvature pinching result for asymptotically conical, locally conformally flat gradient expanding solitons with positive scalar curvature.
@@
Theorem 1.2 is an analogue of Z.-H. Zhang’s result for locally conformally flat gradient shrinking (and steady) solitons [90]. Moreover, by [21, 28], the locally conformally flat assumption can be replaced by the weaker assumption of the vanishing $D$-tensor introduced in [20, 21].
As a consequence, by combining Theorem 1.2 with the work of Cao-Chen [21] and Chen-Wang [41], we have the following application in dimension four.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact, half-conformally flat, asymptotically conical gradient expanding Ricci soliton. Suppose the asymptotic cone has positive scalar curvature. Then, $(M^{4},g,f)$ has positive curvature operator.
Remark 1.3.
By [19, Theorem 5.8], the expanding solitons in Theorem 1.2 and Corollary 1.1 are rotationally symmetric; see also Cao-Yu [28, Corollary 3.4].
Our next two results concern curvature pinching of 4-dimensional asymptotically conical gradient expanding Ricci solitons with either positive isotropic curvature (PIC) or half-positive isotropic curvature (half-PIC).
Recall that, for any oriented 4-manifold $(M^{4},g)$, $2$-forms admit the decomposition $\wedge^{2}(M)=\wedge^{+}(M)\oplus\wedge^{-}(M)$, into self-dual and anti-self-dual $2$-forms. Accordingly, the Riemann curvature operator admits a block decomposition
$Rm=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\frac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\frac{R}{12}I\end{pmatrix}$,
where $W^{\pm}$ denote the self-dual and anti-self-dual part of the Weyl tensor, and $\mathring{Rc}$ the traceless Ricci part. It turns out that $(M^{4},g)$ has PIC (half-PIC) if and only if $A$ and $C$ (AA or $C$) are 2-positive [64]; see Section 2.2 for more details.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with half-positive isotropic curvature (half-weakly PIC). If the asymptotic cone has positive scalar curvature and satisfies either $A\geq 0$ or $C\geq 0$, then $(M^{4},g,f)$ has $A>0$ ($A\geq 0$) or $C>0$ ($C\geq 0$).
@@
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with positive isotropic curvature (weakly PIC). If the asymptotic cone has positive scalar curvature, then the Ricci curvature of $(M^{4},g,f)$ is $2$-positive (22-nonnegative), and $|Rm|\leq 2R$.
@@
Theorems 1.3 and 1.4 are analogues of our previous results for $4$-dimensional complete ancient solutions with half-PIC [26, Proposition 3.1] and PIC [27, Theorem 1.1], respectively. Moreover, both results are valid under some slightly weaker assumptions; see Theorem 4.1 and Theorem 4.2 in Section 4.
By observing the common pattern in the proofs of Theorem 1.3 and Theorem 1.4, we formulate a general method of proof (Lemma 5.1) and apply it to obtain analogues of several additional known generalized Hamilton–Ivey type curvature pinching results for ancient solutions. This includes the following result for $4$-dimensional asymptotically conical gradient expanding Ricci solitons with uniformly positive isotropic curvature (UPIC), as well as several other results as stated in Theorem 5.1.
@@
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact gradient expanding Ricci soliton with uniformly positive isotropic curvature. If $(M^{4},g,f)$ is asymptotic [^3] Here, the notion of asymptotically conical is $C^{2}$-asymptotic in the sense of Chodosh [45, Definition 1.1] which is slightly stronger than that in our Definition 2.1. to a non-flat Euclidean cone, then it is rotationally symmetric.
Remark 1.6.
We note that any $4$-dimensional non-compact gradient expanding Ricci soliton with PIC that is asymptotic to a non-flat Euclidean cone automatically has UPIC. Therefore, one could replace the assumption of UPIC by PIC in both Theorem 1.5 and Corollary 1.2.
@@
Acknowledgements. The first author was partially supported by a Simons Fellowship and a grant from the Simons Foundation. The second author would like to thank Prof. Xiaochun Rong for his continual support and encouragement. Both authors would like to thank Dr. P.-Y. Chan for very helpful information that allowed the removal of the positive scalar curvature assumption on $(M^{n},f,g)$ in Theorem 1.1 and Theorem 1.2 in an earlier version of the paper. We also thank the anonymous referees for their helpful comments.
@@ 2.2. Curvature decomposition and isotropic curvature of four-manifolds
${Rm}=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\tfrac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\tfrac{R}{12}I\end{pmatrix}$, (2.1)
@@
Let $A_{1}\leq A_{2}\leq A_{3}$ and $C_{1}\leq C_{2}\leq C_{3}$ denote the eigenvalues of $A$ and $C$, respectively. It is also well-known that $\operatorname{tr}A=\operatorname{tr}C={R}/{4}$.
@@
It turns out that, in dimension four, these curvature conditions (and their natural extensions, half-PIC or half-WPIC) can be characterized in terms of the $3\times 3$ matrices $A$ and $C$ as follows:
@@
  • half-PIC (half-WPIC) if and only if either $A$ or $C$ is 2-positive (weakly 2-positive), i.e., $A_{1}+A_{2}>0$ ($A_{1}+A_{2}\geq 0$) or $C_{1}+C_{2}>0$ ($C_{1}+C_{2}\geq 0$).
@@ 2.3. Basic differential equations satisfied by curvatures of Ricci solitons
Except for the first identity in Lemma 2.1, the factor 22 differs from [61] due to our normalization of the inner product on $\wedge^{2}(M)$ (see (A.6)). Moreover, the first two equations are valid in all dimensions.
@@ 3. Curvature pinching for asymptotically conical Ricci expanders with vanishing Weyl tensor
This section is devoted to the proof of Theorem 1.1 and Theorem 1.2. Note that both theorems follow from the following result.
@@
First of all, since the asymptotic cone of $(M^{n},g,f)$ has positive scalar curvature, it follows from the strong maximum principle and [33, Theorem 1.6] that $(M^{n},g,f)$ itself has positive scalar curvature $R>0$. This fact will be used in the proofs of Theorem 3.1 and Theorem 1.1.
Next, we shall need the following differential inequality for the ratio between the smallest Ricci eigenvalue $\lambda_{1}$ and the scalar curvature $R$.
@@
Let $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$ be the Ricci eigenvalues. Fix any point $p_{0}\in M$, and choose a unit vector field $e_{1}$ such that $Rc(e_{1})=\lambda_{1}e_{1}$ at $p_{0}$. Extend $e_{1}$ by parallel transport along geodesics emanating from $p_{0}$. Clearly, $\lambda_{1}\leq Rc(e_{1},e_{1})$ with equality at $p_{0}$. Evaluating at $p_{0}$, we obtain, in the barrier sense (cf. Calabi [14]),
@@
where in the last equality, we have used [55, (2.8)] (see also [81, Lemma 2.5]).
@@
Finally, it is clear from the above expression that $h_{1}\leq 0$. ∎
Now, we divide the proof of Theorem 3.1 into three parts.
Part I: $(M^{n},g,f)$ has nonnegative Ricci curvature $Rc\geq 0$.
@@
Case 1: Interior infimum. Suppose the negative infimum $\epsilon_{1}$ is attained at some point $p_{0}\in M$. By Lemma 3.1 and Calabi’s barrier strong maximum principle (Lemma 2.2), the ratio $\lambda_{1}/R$ must be constant on some bounded connected domain $p_{0}\in\Omega\subset M^{-}$. In particular, $h_{1}\equiv 0$ on $\Omega$ which forces
@@
On the other hand, since the expanding Ricci soliton satisfies $W\equiv 0$, its asymptotic cone $\mathcal{C}$ also has vanishing Weyl tensor. By Lemma A.1 (for $n=3$) and Lemma A.2 (for $n\geq 4$), $\mathcal{C}$ has nonnegative Ricci curvature, whose smallest eigenvalue $\tilde{\lambda}_{1}\equiv 0$ occurs in the radial direction. Hence
@@
Part II: $(M^{n},g,f)$ has nonnegative curvature operator $Rm\geq 0$.
First of all, when the Weyl tensor $W$ vanishes, we may express the smallest sectional curvature as (see, e.g., [90, Proposition 3.1])
@@
We also note that, for any manifold with vanishing Weyl tensor, $\mu$ also coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [90, Proposition 3.1(i)]). In particular, when $W=0$, nonnegative sectional curvature is equivalent to nonnegative curvature operator $Rm\geq 0$.
@@
Evaluating at $p_{0}$ in the barrier sense, using (3.6) and [55, (2.8)] (see also [81, Lemma 2.5]), we obtain
@@
$\begin{split}\bigl((n-2)\lambda_{2}-R\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}&=-\Bigg(\lambda_{1}+\lambda_{2}+\sum_{j=3}^{n}(\lambda_{j}-\lambda_{2})\Bigg)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &=-(\lambda_{1}+\lambda_{2})\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &\quad-\Biggl(\sum_{k=3}^{n}(\lambda_{k}-\lambda_{2})\Biggr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}.\end{split}$ (3.8)
@@
Now, we are ready to complete the proof that $(M^{n},g,f)$ has $Rm\geq 0$.
Proof of $Rm\geq 0$.
@@
Case 1: Interior infimum. Suppose the negative infimum $\epsilon_{0}$ is attained at some point $p_{0}\in M$. By Part I of the proof above, we know that the Ricci curvature is nonnegative. Then, by Lemma 3.2, we have